OpenStudy (anonymous):

multiply (x+1)(x−4)(x−√3) into standard polynomial form

OpenStudy (anonymous):

I got x3−3x2−4x−x23√+3x3√+43√ but i graphed it and the zeros were not right

OpenStudy (anonymous):

or do imaginary numbers come into play?

OpenStudy (chriss):

x^3+sqrt(3) x^2+5 x^2+5 sqrt(3) x+4 x+4 sqrt(3)

OpenStudy (anonymous):

\[x ^{3}-3x ^{2}-x ^{2}\sqrt{3}+3x \sqrt{3}+4\sqrt{3}-4x\]

myininaya (myininaya):

\[(x-a)(x-b)(x-c)=(x^2-bx-ax+ab)(x-c)\] \[=x^3-cx^2-bx^2+bcx-ax^2+acx+abx-abc\] \[=x^3-(a+b+c)x^2+(ab+ac+bc)x-abc\]

OpenStudy (anonymous):

When I put yours in Chris the zeros came out right on the calculator. May I ask how you came to the answer you got? when i multiply the square root is where i get messed up.

myininaya (myininaya):

\[a+b+c=-1+4+\sqrt{3}=3+\sqrt{3} \] \[ab+ac+bc=-1(4)+(-1)(\sqrt{3})+4\sqrt{3}=-4-1\sqrt{3}+4\sqrt{3}=-4+\sqrt{3} \] \[abc=-1(4)\sqrt{3}=-4\sqrt{3}\]

OpenStudy (anonymous):

i am going to make a bet, which is that whoever gave you this problem meant "integer coefficients" not "real coefficients" and so you were not supposed to do all this. that is just a guess though

OpenStudy (anonymous):

when you originally posted i though you should use \[(x+1)(x-4)(x-\sqrt{3})(x+\sqrt{3})\] and that will give you integer coefficients. of course i could be wrong because it did say "real" not "integer"

OpenStudy (anonymous):

The problem said write a polynomial function using the given zeros

OpenStudy (anonymous):

and those were -1,4, sqt3

OpenStudy (anonymous):

i put those in the equation i put above

myininaya (myininaya):

he's wondering about the coefficients though does it say real or integer coefficients?

myininaya (myininaya):

or does it not say anything?

OpenStudy (anonymous):

i don't have it in front of me but I'm pretty sure it did not say

OpenStudy (anonymous):

What chris posted though worked and i got the zeros when i graphed it

OpenStudy (chriss):

tbh, I popped it into Wolfram alpha to get that

myininaya (myininaya):

ok i have looked at your posting before this and it said find a polynomial of least degree having these zeros it does say least degree if so then you are correct in saying (x+1)(x-4)(x-sq3)

OpenStudy (anonymous):

But then you are suppose to put it into polynomial form.

myininaya (myininaya):

standard form? i wrote this for you in standard form above

OpenStudy (anonymous):

\[x^3-\sqrt{3} x^2-3 x^2+3 \sqrt{3} x-4 x+4 \sqrt{3}\] is one answer http://www.wolframalpha.com/input/?i=%28x%2B1%29%28x-4%29%28x-sqrt%283%29%29